# Beyond time-homogeneity for continuous-time multistate Markov models

### v2

@@ 1 $1-p_{1}$ $p_{1}$ 0 0
For the simulation study, we compare the four approaches to fitting the HMM as outlined in Section 3.2. Furthermore, the matrix-exponential-based solutions (C) and (D) are fitted – for each of the 100 synthetic data sets – for each choice of $d\in\{1/6,1,2\}$, as in (4), representing time-homogeneity at the resolution of bi-monthly, yearly, and bi-yearly, respectively. The solutions (C) and (D) become computationally prohibitive at the monthly resolution, i.e., $d=1/12$. A summary of the results is displayed in Figures 2 and 3; violin plots of the posterior means or MLEs for the baseline coefficient parameters for each transition rate are shown, along with the coverage of .95 level credible or confidence intervals. Note that solutions (C) and (D) represent the model fit when it is misspecified because the data are simulated with the rate matrix updating continuously as a function of time, not piecewise-constant with respect to time. Thus, solutions (A) and (B) are fit using a correctly specified model. The corresponding violin plots for all HMM parameters are available in the supplementary material.
@@
Figure 2: For approaches (A) and (C), presented are violin plots of the posterior means corresponding to the simulation study of 100 synthetic data sets. The violin plots associated with approaches (B) and (D) are of the MLEs. The baseline coefficient parameters for the transition rates are as specified in equation (7). The numbers below each label report coverages of .95 level credible or confidence intervals.
@@
Figure 3: For approaches (A) and (C), presented are violin plots of the posterior means corresponding to the simulation study of 100 synthetic data sets. The violin plots associated with approaches (B) and (D) are of the MLEs. The baseline coefficient parameters for the transition rates are as specified in equation (7). The numbers below each label report coverages of .95 level credible or confidence intervals.
@@
Consistent with the intuition provided in Section 2.1, it is evident from Figures 2 and 3 that substantial estimation bias for the baseline rate parameters (i.e., $\beta_{0,j}$ for $j\in\{1,\dots,5\}$) is exhibited by the HMMs fitted with the time-homogeneous approximation (4), for each choice of $d\in\{1/6,1,2\}$, whereas the coefficient estimates based on the numerical ODE solver approaches to fitting the time-inhomogeneous HMM (3) are well-concentrated on the true coefficient values. Note that despite approach (B) leading to unbiased estimates, the coverage probabilities for $\beta_{0,1}$, $\beta_{0,4}$, and $\beta_{0,5}$ are markedly less than the nominal level of .95, in contrast to those of (A). Moreover, less time is required for computing the likelihood function based on the numerical ODE solution than is required for the matrix-exponential-based computations for $d=1/6$, as shown in Table 1. There are observed, however, reduced computation times for the matrix-exponential-based solutions for $d\in\{1,2\}$, but these come at the cost of more substantial estimation bias.
@@ 3.4.1 Cardiac allograft vasculopathy study data
As described in the simulation study in Section 3.3, the continuously time-evolving transition rates were constructed as a feature of the synthetic or artificial CAV data (in order to assess the estimation performance observed when fitting the HMMs to data that are actually generated from a Markov process with this feature). Instead, in this section, we model the real CAV data set in the same manner as Section 3.3 and compare the four approaches (A), (B), (C), and (D), except now, time-evolving transition rates may not be a feature exhibited by these data, but in that case the slope coefficients for $t$ (i.e., $\beta_{1,j}$ for $j\in\{1,\dots,5\}$) would be estimated as zero and the four approaches would largely yield the same inference on the baseline rates. Thus, any observed difference in the estimated transition rates is likely attributable to time-evolving rates exhibited by the data, and it would be rather contrived to assume that the underlying rates of CAV progression actually evolve in a piecewise-constant fashion – let-alone to identify the correct resolution of the discontinuity points. Hence, any discrepancies in the resulting inference from the four approaches is indicative that future practitioners applying continuous-time, time-inhomogeneous Markov models should make a conscious and justified decision for which method of evaluating the transition probability matrix is used.
@@
In any case, the .95 level credible or confidence sets for the baseline rate parameters – fitted by all four approaches and for each $d\in\{1/6,1,2\}$ – are presented in Figure 4. Differences are observed in the intervals [across the estimation approaches] most prominently between the numerical ODE solver approaches versus the piecewise-homogeneous approximations for lower resolutions, $d$. Comparing numerical approaches (A) and (B), the confidence intervals of (B) for the baseline parameters tend to be more narrow, with the exception of $\beta_{0,2}$. This wide confidence set phenomenon for approach (B) is also seen in the confidence sets for the initial state parameters (provided in the supplementary material). For completeness, Table 2 outlines the computation speed for the likelihood evaluations using the different approaches. The same trend as in the simulation study of Section 3.3 manifests. Finally, there is one special case (that we are aware of) in which the piecewise-homogeneous approach (for some $d>0$) may be a reasonable approximation to the numerical ODE solution, for fitting a time-inhomogeneous HMM with a continuously evolving rate matrix; this special case is discussed in the next section.
@@
Figure 4: For each parameter (i.e. within the dotted vertical lines), the vertical intervals ordered left to right correspond to the “Approach” ordered top to bottom. For approaches (A) and (C), highest posterior density credible intervals are reported for the .95 level, for the fitted baseline coefficients, $\beta_{0,j}$, as in equation (7), for all transitions. For approaches (B) and (D), .95 level confidence intervals based on the Hessian are reported, as given by the nhm (Titman, 2019) and msm packages (Jackson, 2011), respectively.
approach (A) (B): $d=1/6$ (B): $d=1$ (B): $d=2$
@@ 4 Concluding remarks
Continuous-time multistate Markov models (HMMs or otherwise) are widely used in data analysis applications involving processes that evolve over time, and generally speaking it is artificial to assume that a naturally occurring process would exhibit rates of evolution that change in an arbitrarily regular piecewise-constant fashion. Although this assumption motivates a pragmatic approximation for fitting a time-inhomogeneous, continuous-time Markov model, we demonstrate that: such an approximation may lead to non-negligible biases in estimation; there exist exact numerical solutions that are also computationally more efficient; and the approximation is unnecessary – the computational tools to implement the exact numerical solutions are already readily and freely available.
5 Appendix: Proofs

### v3

@@ 1 $1-p_{1}$ $p_{1}$ 0 0
For the simulation study, we compare the five approaches to fitting the HMM as outlined in Section 3.2. Furthermore, the matrix-exponential-based solutions (C) and (D) are fitted – for each of the 100 synthetic data sets – for each choice of $d\in\{1/6,1,2\}$, as in (5), representing time-homogeneity at the resolution of bi-monthly, yearly, and bi-yearly, respectively. The solutions (C) and (D) become computationally prohibitive at the monthly resolution, i.e., $d=1/12$. A summary of the results is displayed in Figures 2 and 3; violin plots of the posterior means or MLEs for the baseline coefficient parameters for each transition rate are shown, along with the coverage of .95 level credible or confidence intervals. Note that solutions (C) and (D) represent the model fit when it is misspecified because the data are simulated with the rate matrix updating continuously as a function of time, not piecewise-constant with respect to time. Next, solution (E) represents a model fit when it is misspecified for the following three reasons: states are observed with error, not all transitions between states are necessarily observed, and observation times do not necessarily correspond to transition times. Only solutions (A) and (B) are fit using a correctly specified model. Violin plots for all HMM parameters are in the supplemental materials.
@@
Figure 2: For approaches (A) and (C), presented are violin plots of the posterior means corresponding to the simulation study of 100 synthetic data sets. The violin plots associated with approaches (B) and (D) are of the MLEs. The violin plots for approach (E) are from the empirical estimation process outlined in Section 5.2. The baseline coefficient parameters for the transition rates are as specified in equation (8). Where available, the numbers below each label report coverages of .95 level credible or confidence intervals.
@@
Figure 3: For approaches (A) and (C), presented are violin plots of the posterior means corresponding to the simulation study of 100 synthetic data sets. The violin plots associated with approaches (B) and (D) are of the MLEs. The violin plots for approach (E) are from the empirical estimation process outlined in Section 5.2. The baseline coefficient parameters for the transition rates are as specified in equation (8). Where available, the numbers below each label report coverages of .95 level credible or confidence intervals.
@@
Consistent with the intuition provided in Section 2.1, it is evident from Figures 2 and 3 that substantial estimation bias for the baseline rate parameters (i.e., $\beta_{0,j}$ for $j\in\{1,\dots,5\}$) is exhibited by the HMMs fitted with the time-homogeneous approximation (5), for each choice of $d\in\{1/6,1,2\}$, whereas the coefficient estimates based on the numerical ODE solver approaches to fitting the time-inhomogeneous HMM (4) are well-concentrated on the true coefficient values. Additionally, a similar bias is observed with regards to approach (E). Note that despite approach (B) leading to unbiased estimates, the coverage probabilities for $\beta_{0,1}$, $\beta_{0,4}$, and $\beta_{0,5}$ are markedly less than the nominal level of .95, in contrast to those of (A). Moreover, less time is required for computing the likelihood function based on the numerical ODE solution than is required for the matrix-exponential-based computations for $d=1/6$, as shown in Table 1. There are observed, however, reduced computation times for the matrix-exponential-based solutions for $d\in\{1,2\}$, but these come at the cost of more substantial estimation bias.
@@ 3.4.1 Cardiac allograft vasculopathy study data
As described in the simulation study in Section 3.3, the continuously time-evolving transition rates were constructed as a feature of the synthetic or artificial CAV data (in order to assess the estimation performance observed when fitting the HMMs to data that are actually generated from a Markov process with this feature). Instead, in this section, we model the real CAV data set in the same manner as Section 3.3 and compare the five approaches (A), (B), (C), (D), and (E) except now, time-evolving transition rates may not be a feature exhibited by these data, but in that case the slope coefficients for $t$ (i.e., $\beta_{1,j}$ for $j\in\{1,\dots,5\}$) would be estimated as zero and the five approaches should largely yield the same inference on the baseline rates. Thus, any observed difference in the estimated transition rates is likely attributable to time-evolving rates being exhibited by the data, and it would be rather contrived to assume that the underlying rates of CAV progression actually evolve in a piecewise-constant fashion – let-alone to identify the correct resolution of the discontinuity points. Similarly, there is no reason to believe that the progression of CAV occurs precisely at the observed time points. Hence, any discrepancies in the resulting inference from the five approaches is indicative that future practitioners applying continuous-time, time-inhomogeneous Markov models should make a conscious and justified decision for which method of evaluating the transition probability matrix is used.
@@
In any case, the .95 level credible or confidence sets for the baseline rate parameters – fitted by all five approaches and for each $d\in\{1/6,1,2\}$ – are presented in Figure 4. Differences are observed in the intervals [across the estimation approaches] most prominently between the numerical ODE solver approaches versus the piecewise-homogeneous approximations for lower resolutions, $d$, and approach (E). Comparing numerical approaches (A) and (B), the confidence intervals of (B) for the baseline parameters tend to be narrower, with the exception of $\beta_{0,2}$. This wide confidence set phenomenon for approach (B) is also seen in the confidence sets for the initial state parameters (provided in the supplemental materials). For completeness, Table 2 outlines the computation speed for the likelihood evaluations using the different approaches. The same trend as in the simulation study of Section 3.3 manifests. Finally, there is one special case (that we are aware of) in which the piecewise-homogeneous approach (for some $d>0$) may be a reasonable approximation to the numerical ODE solution, for fitting a time-inhomogeneous HMM with a continuously evolving rate matrix; this special case is discussed in the next section.
@@
Figure 4: For each parameter (i.e., within the dotted vertical lines), the vertical intervals ordered left to right correspond to the “Approach” ordered top to bottom. For approaches (A) and (C), highest posterior density credible intervals are reported for the .95 level, for the fitted baseline coefficients, $\beta_{0,j}$, as in equation (8), for all transitions. For approaches (B) and (D), .95 level confidence intervals based on the Hessian are reported, as given by the nhm (Titman, 2019) and msm packages (Jackson, 2011), respectively. Due to the lack of availability of an approach for constructing a confidence interval for approach (E), only point estimates are provided.
approach (A) (C): $d=1/6$ (C): $d=1$ (C): $d=2$
@@ 4 Concluding remarks
Continuous-time multistate Markov models (HMMs or otherwise) are widely used in data analysis applications involving processes that evolve over time, and generally speaking it is artificial to assume that a naturally occurring process would either exhibit rates of evolution that change in an arbitrarily regular piecewise-constant fashion, or that all observation times correspond to state transition times. Although these assumptions motivate pragmatic approximations for fitting a time-inhomogeneous, continuous-time Markov model, we demonstrate that such approximations may lead to non-negligible biases in estimation. Furthermore, we demonstrate that there exist exact numerical-integration solutions. Moreover, these solutions are computationally more efficient than the approximations, and the computational tools to implement them are readily and freely available. Given the comprehensive list of approaches tested in this analysis, we find that the numerical-integration approaches, based on either Bayesian (approach (A)) or Frequentist (approach (B)) computations, lead to the most consistent estimation of the standard parametric structures that we consider in the simulation study in Section 3.3 and real data analyses in Section 3.4. Lastly, we note that the numerical integration can be implemented via the Riemann integral (by way of deSolve and/or nhm) or the product integral (by way of prodint).
5 Appendix
5.1 Proofs
@@
5.2 Transition rate parameter estimation via the Nelson-Aalen estimator
Recall that the transition rate matrix, ${\bf\it Q}(t)$, from Sections 3.3 and 3.4.1 has the form
where each $q_{j}$, $j=1,2,\ldots,5$, has the following form
$q_{j}=q_{j}(t\mid\text{sex})=\exp(\beta_{0,j}+\beta_{1,j}\cdot t+\beta_{2,j}\cdot\text{sex}),$
with $t\geq 0$ and $\text{sex}\in\{0,1\}$. The Nelson-Aalen estimator estimates the cumulative transition rate function (Aalen & Johansen, 1978; Borgan, 1997). Let $A_{j}(t\mid\text{sex})$ be the cumulative transition rate function defined as
[…]
and let $\hat{A}_{j}(t\mid\text{sex})$ denote the Nelson-Aalen estimator. The aalen_johansen function provides $\hat{A}_{j}(t\mid\text{sex})$, for each observed transition time, $t\geq 0$, and for $j=1,2,\ldots,5$. Then, in order to retrieve estimates for $\beta_{0,j}$, $\beta_{1,j}$, and $\beta_{2,j}$, $\forall j$, define
[…]
$=\frac{\exp(\beta_{0,j}+\beta_{1,j}\cdot t+\beta_{2,j}\cdot\text{sex})-\exp(\beta_{0,j}+\beta_{2,j}\cdot\text{sex})}{\beta_{1,j}},$
and use the $\ell_{1}$ loss to formulate the parameter estimates as
$\widehat{{\bf\it\beta}}_{j}^{T}=\matrixquantity(\hat{\beta}_{0,j}\\ \hat{\beta}_{1,j}\\ \hat{\beta}_{2,j})=\underset{\beta_{0,j},\beta_{1,j},\beta_{2,j}}{\textup{argmin}}\sum_{t}\quantity|g(\beta_{0,j},\beta_{1,j},\beta_{2,j}\mid t,\text{sex})-\hat{A}_{j}(t\mid\text{sex})|.$